OpenStudy (mendicant_bias):

Having a problem with solving an IVP using Laplace involving Heavisides, work so far posted in a moment.

OpenStudy (mendicant_bias):

Use the Laplace Transform to solve the given Initial Value Problem. \[y' +y=f(t), \ \ \ y(0)=0 \]where \[f(t)= \begin{cases} 0, \ \ \ 0 \leq t<1 \\ 5, \ \ \ t \ge 1 \end{cases} \]

OpenStudy (mendicant_bias):

First, putting the whole thing in the form of a bunch of linear combinations (turning that f(t) into a bunch of heaviside-multiplied functions)

OpenStudy (mendicant_bias):

\[f(t)=5\mathcal{U}(t-1)\]

OpenStudy (jhannybean):

@perl @Concentrationalizing

OpenStudy (mendicant_bias):

\[y'+y=5 \ \mathcal{U}(t-1)\]

OpenStudy (mendicant_bias):

Taking the laplace of all of that:\[[sY(s)-y(0)]+Y(s)=\frac{5e^{-s}}{s}\]

OpenStudy (mendicant_bias):

\[Y(s)(s+1)=\frac{5e^{-s}}{s}\]

OpenStudy (mendicant_bias):

\[Y(s)=\frac{5e^{-s}}{s(s+1)}\]

OpenStudy (mendicant_bias):

\[\mathcal{L^{-1}}(Y(s))=5 \mathcal{L^{-1}}\left\{\vphantom{} \frac{e^{-s}}{s(s+1)} \right\}\]

OpenStudy (mendicant_bias):

By Partial Fraction Decomposition,

OpenStudy (mendicant_bias):

\[\frac{1}{s(s+1)}=\frac{1}{s}-\frac{1}{s+1}\]

OpenStudy (mendicant_bias):

By making the identification that,\[a=-1, \ \ \ f(t)=\Bigg[\frac{1}{s}-\frac{1}{s+1}\Bigg]\]We can figure out that our solution is going to be of the form,

OpenStudy (mendicant_bias):

\[f(t-1)\mathcal{U}(t-1)\]

OpenStudy (mendicant_bias):

\[\mathcal{L^{-1}}\left\{\Bigg[\frac{1}{s}-\frac{1}{s+1}\Bigg]\right\}=1-e^{-s}\]

OpenStudy (mendicant_bias):

Multiplying that by the transformed Unit Step and accounting for the shift, we have:\[5 \ \mathcal{U}(t-1)(1-e^{-s-1})\]?

OpenStudy (mendicant_bias):

Whoops, those last two variables supposed to be t's, not s's.

OpenStudy (mendicant_bias):

(Lol solved my own problem in the process, nvm, don't need help; I forgot about something early on that messed me up before the transform).

OpenStudy (mendicant_bias):

And the exponent is supposed to be,

OpenStudy (mendicant_bias):

\[e^{-(t-1)}\]

ganeshie8 (ganeshie8):

\[Y(s)=5e^{-s}\left( \dfrac{1}{s(s+1)}\right) \implies y(t) = 5\mathcal{U}(t-1)(1-e^{-(t-1)})\] like this ?

OpenStudy (mendicant_bias):

Yeah!

ganeshie8 (ganeshie8):

looks good to me